OpenStudy (anonymous):

Which is an expression to find the speed in m/s? d = 100 m, t = 20 s A. 100 - 20 B. 100 + 20 C. 100 / 20 D. 100 x 20

OpenStudy (anonymous):

@flasht

OpenStudy (anonymous):

You know that d = v * t?

OpenStudy (anonymous):

d

OpenStudy (anonymous):

@flasht Please don't give direct answers. :)

OpenStudy (anonymous):

@aleroth that makes no sense to me

OpenStudy (anonymous):

Okay. "d = v * t" can be interpreted as "distance = speed * time."

OpenStudy (anonymous):

If you just remember that one, you will be able to solve every equation that involves these 3 variables.

OpenStudy (anonymous):

ahh distance equals speed times time?

OpenStudy (anonymous):

Indeed

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

In this case you want speed, so you divide both sides with time to get speed alone. speed = distance/time

OpenStudy (anonymous):

so it is D?

OpenStudy (anonymous):

Not quite. You have: d = 100 m t = 20 s So: s = (100 m) / (20 s) = (5 m /s)

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

No, you need to DIVIDE distance with time to get speed, which is what you are after. :)

OpenStudy (anonymous):

sooo C? @aleroth

OpenStudy (anonymous):

Yeah, it's C. I hope you understand why.

OpenStudy (anonymous):

i do thank you

OpenStudy (anonymous):

Remember these: d = vt v = d/t t = d/v